OpenStudy (anonymous):

1.Write the equation of a line parallel to the line 2x + 3y = 5 and passes through the point (9, -3) 2. Write the equation of a line in slope-intercept form that has a slope of -2 and passes through the point (2, -8). A. y = -2x - 4 B. y = -2x - 8 C. y = -2x + 18 D. y = -2x – 12 3. Write an equation of a line in point-slope form that has a slope of -2 and passes through (5, -1). A. y + 1 = -2(x – 5) B. y – 1 = -2(x – 5) C. y – 5 = -2(x + 1) D. y -5 = -2(x – 1)

OpenStudy (john_es):

1) If you want a parallel line to ax+by=c, it must be of the form, ax+by=n, where you must find the value of n using the point they give you. Do you understand it?

OpenStudy (anonymous):

not really.

OpenStudy (john_es):

For example, a line parallel to the line 2x+3y=5 is, 2x+3y=8. Do you see what they have in common?

OpenStudy (anonymous):

yeah the 2x+3y

OpenStudy (john_es):

Perfect, so in your problem the line must be that way: 2x+3y=n, but you don't know the value of n. Ok?

OpenStudy (anonymous):

ok

OpenStudy (john_es):

Well, we must use the point the give us and substitute in our equation, so (x,y)=(9,-3), that is, x=9 and y=-3, \[2x+3y=n\Rightarrow 2\cdot 9+3\cdot (-3)=n\]What value do you obtain for n?

OpenStudy (anonymous):

9?

OpenStudy (anonymous):

whit nvm lemme do that again

OpenStudy (john_es):

Perfect, so the solution is 2x+3y=9.

OpenStudy (anonymous):

oh so it is 9?

OpenStudy (john_es):

Yes, the solution to the first problem is 2x+3y=9. :)

OpenStudy (anonymous):

okay(: thanks. can you help with the other 2 now?

OpenStudy (john_es):

Of course. For the second question you should remember the point-slope form of the line, \[y=mx+n\]where m is the slope, and n the y intercept. With the data of the problem and the point-slope equation, could you say me what is the correct answer?

OpenStudy (anonymous):

would it be c?

OpenStudy (anonymous):

-2x+18?

OpenStudy (john_es):

As you see, all lines satisfy the condition: slope=-2. So our equation should be of the form, \[y=-2x+n\]But we neew to find n. Again you must substitue x and y with the values they give you, x=2 and y=-8 (but it is not c).

OpenStudy (john_es):

Other form to check it is to do x=2 in each solution, and see what of the equations gives you y=-8.

OpenStudy (anonymous):

4

OpenStudy (john_es):

with a minus sign, so it should be A, y=-2x-4, because when you do x=2, you obtain y=-8. Ok?

OpenStudy (anonymous):

okay so it's A?

OpenStudy (john_es):

Yes, by the reason I gave you.

OpenStudy (anonymous):

okay(: & now can you help me with the last one?

OpenStudy (john_es):

The third problems it should be easier. The point intercept equation has the form, \[y-y_0=m(x-x_0)\]In the problem says you slope=m=-2, and the point is (x0,y0)=(5,-1). Could you select the correct answer with this data?

OpenStudy (anonymous):

my guess would be either C or D

OpenStudy (john_es):

I give you a hint. \[y-(-1)=-2(x-5)\]:)

OpenStudy (anonymous):

a?

OpenStudy (john_es):

Perfect. I hope it helps you (but I think you must rework it a little).